OpenStudy (anonymous):

Evaluate the triple integral H of (x^2 + y^2 + z^2)^2 d, where H is the solid hemisphere x^2 + y^2 + z^2 less than or equal to 9 with z greater than or equal to 0

OpenStudy (amistre64):

what are your thoughts on it?

OpenStudy (anonymous):

well I understand how to take a triple integral, but I always get confused on the limits. Would I put it in terms of the spherical coordinate system? such as \[\int\limits_{0}^{\pi} \int\limits_{0}^{2\pi} \int\limits_{0}^{3} (e^2 \sin \phi)^2 de d \Theta d \phi\]

OpenStudy (anonymous):

I would say that's almost right. Check your upper bound on \(\phi\). What you have suggests that we also consider the half of the sphere below the \(x\)-\(y\) plane.

OpenStudy (anonymous):

\[z \ge0\], so it wouldn't be below the xy-plane right?

OpenStudy (anonymous):

That's right. However, \(z=e\cos\phi\). But if \(\dfrac{\pi}{2}<\phi<\pi\), you have \(-1<\cos\phi<0\), which would mean \(z\) is negative. This means you have to use \(0\le \phi\le\dfrac{\pi}{2}\) as your interval.

OpenStudy (anonymous):

ohh okay that makes a lot of sense, so if I change phi the rest of the set up is correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

is the answer \[\frac{ 243\pi^2 }{ 10 }\]

OpenStudy (anonymous):

Yep I'm getting the same solution.